OpenStudy (anonymous):

What is the value of log 5? 25 And could you tell me how you got the answer please.

OpenStudy (anonymous):

log 25=2log5

OpenStudy (anonymous):

log 5=0.698970004

OpenStudy (anonymous):

I think they meant \[\log_{25} (5)\]

OpenStudy (anonymous):

y= log{25} 5 25^y = 5 5^(2y) = 5 2y=1 y=(1/2)

OpenStudy (anonymous):

answer = (1/2)

OpenStudy (anonymous):

Thanks